HOW MUCH (ch4) DOES IT TAKE TO GENERATE 1GWH OF ELECTRICITY


1
Expert's answer
2021-05-18T11:09:20-0400

1 kg of methane (CH4) produces 55.50 MJ of energy. Assuming that 100% of heat of combustion is used to produce electricity, the mass required to generate 1 GWh of energy can be calculated by dividing the value of electrical energy and heat of combustion expressed in the same units:


"E_\text{electrical}=1·10^9·3600=3.6·10^{12}\text{ J}."

The mass of methane:


"m=\frac Eq=\frac{3.6·10^{12}}{55·10^6}=720·10^3\text{ kg}."